OpenStudy (anonymous):

sqrt(c^5) * sqrt(c^6)^1/2

OpenStudy (anonymous):

\[\large \sqrt{c^6}=c^3\] is a start

OpenStudy (anonymous):

then \[\large \sqrt{c^5}\sqrt{c^3}=\sqrt{c^8}=c^4\]